# Alleviating Non-identifiability: a High-fidelity Calibration Objective for Financial Market Simulation with Multivariate Time Series Data

### v3

@@ 4.2 Results and Analysis
Calibration to the synthetic data. Note that, although the non-identifiable set of any $F_{i}$ is verified to be smaller than that of any $F_{j}$, $i>j$, it does not deterministically lead to better calibration performance of $F_{i}$ than $F_{j}$. Because $F$ mainly decides the parameter space, while the calibration performance is also jointly influenced by the optimization algorithms. Given the calibration problem is black-box with multiple local optima, it is non-trivial for the optimization algorithms to find the global optimum even the parameter space is modeled fully identifiable.
The performance between each simulated data and its corresponding synthetic data is measured by the indicators based on both MSE and W, respectively. In Figure 6, each (blue or yellow) point displays the average performance on these 10 synthetic data. It shows that by using more features in $F$, the performance of calibration generally improves as the distance values decrease.
Furthermore, it is interesting to see how the simulation model calibrated by all features ($F_{6}$) compares to the ones calibrated to each single feature $f_{i},i=1,...,6$, instead of in an incremental manner of $F_{i},i=1,...,5$. As listed in Table 1, by calibrating with $F_{6}$, the obtained parameter enjoys the best simulation performance, i.e., the smallest averaged distances on 7 out of the 10 instances, and top 3 averaged distances on the other 3 instances. Besides, the friedman test is applied to show that $F_{6}$ achieves the best average performance and the advantage is statistically significant. This implies that the proposed theory and new objective function are applicable to various features and not restricted by the feature with more information of the social system, e.g., the mid-price.
@@
#rank 3.20 3.60 3.80 4.50 5.30 3.80 1.57
Table 1: Comparisons of calibration performance between 6 univariate data and $F_{6}$. Note that the performance is at the $1E10^{-2}$ level.
@@
Then, we illustrate the time series of 6 features to show how the simulated data resembles the synthetic data in Figure 7. Specifically, we calibrate the model by $F_{6}$ and $f_{x}$ on the data generated based on the recommended parameter setting $\omega^{*}=[0.025,100,10,0.001,0.15,0.025]$  [11], where $f_{x}$ indicates any single feature who helps to calibrate the model and generate the best simulated data. And the simulated data calibrated by $F_{6}$ resembles the target data on all 6 features very closely. To our best knowledge, this is the first ABM-agnostic method of successfully calibrating the high-frequency market time series at 1 second level and simulating 6 features with high-fidelity. In comparison, the simulated data of $f_{x}$ performs much poorer.
Calibration to the real data. At last, it is verified that the proposed method can also calibrate the real market data at a 3-second level. A 1-hour time series of 000001.sz is calibrated using $F_{1}$ and $F_{3}$. The simulated data generated by the corresponding best calibrated models is depicted in Figure 8, together with the real observed data. Note that, most of the FMS simulators (including the adopted PGPS model) model the order volume in a coarse-grained manner, i,e., simply setting all the orders with the identical volume of 100. Hence, only the mid-price time series is shown in a time step of 1 minute. It can be seen that the simulated mid-price of $F_{3}$ is more similar to the real observed data than $F_{1}$ by an improvement of 27.5% in MSE. This confirms that the proposed method indeed can calibrate the real market data with multiple time series, thus suggesting a new effective calibration objective function for financial market simulation.
@@
5 Conclusions
@@
This paper studies the non-identifiability issue of social simulation. To our best knowledge, this is the first work that the non-identifiability is formally defined and theoretically alleviated. It is defined as the probability of randomly sampling in the non-identifiable set of the parameter space. It is rigorously analyzed that this probability can be reduced with more time series features in an exponential manner. Based on the analysis, a new objective function is proposed to effectively aggregate multiple features with a maximization function over existing discrepancy metrics. Extensive empirical studies have been conducted on 10 synthetic data and 1 real data of the FMS tasks. It has been successfully verified in the simulations that, with multiple features, not only the non-identifiable set can be minimized but the simulation fidelity is largely improved.

### v7

@@ IV-B Results on Alleviation of Non-identifiability
Figure 9: The average Pearson correlation coefficient of the selected features among 10 simulated data.
Building on this, an additional experiment investigates the impact of feature dependencies on the alleviation. The theoretical proof in Appendix A-C suggests that linearly dependent features do not contribute to the alleviation. To validate this, six new test features are constructed based on the original six, where $f_{1}^{\prime}=f_{1},f_{2}^{\prime}=2f_{1},f_{3}^{\prime}=-2f_{1},f_{4}^{\prime}=f_{5},f_{5}^{\prime}=f_{6},f_{6}^{\prime}=f_{2}$. The corresponding new objectives, $F_{i}^{\prime}$, are constructed in the same manner as the original $F_{i}$. These features include both linearly dependent ones ($f_{1}^{\prime},f_{2}^{\prime},f_{3}^{\prime}$) and non-linearly related ones ($f_{4}^{\prime},f_{5}^{\prime},f_{6}^{\prime}$), enabling a direct evaluation of their respective effects on NIP reduction. As illustrated in Fig. 7, adding linearly dependent features does not reduce non-identifiability, aligning with the theoretical findings in Appendix A-C. In contrast, introducing non-linearly related features leads to a significant reduction, confirming that independent features impose additional constraints on the non-identifiability set.
While adding more independent features continues for alleviation, the marginal benefit diminishes as more features are included. This effect arises because the initial features already contribute substantially to the reduction, leaving less room for further alleviation. As a result, even if newly added features are entirely independent or possess highly distinctive and informative properties, their impact on further shrinking the non-identifiability set is inherently limited. Mathematically, this corresponds to their associated $\beta$ values approaching 1, reflecting the decreasing marginal effect of additional features rather than an indication of feature redundancy.
To further assess the effectiveness of the originally selected six features, the average Pearson correlation coefficient among ten target datasets is computed, as shown in Fig. 9. The results indicate low pairwise correlations among the selected features, supporting their effectiveness in reducing NIP and confirming their direct applicability in calibration.
TABLE III: Comparisons of calibration performance between 6 univariate data and $F_{6}$.*
@@
#rank $3.20^{+}$ $4.10^{+}$ $4.00^{+}$ $5.10^{+}$ $5.60^{+}$ $4.50^{+}$ 1.50
    Performance is at the $10^{-2}$ level. $+$ denotes significance under Bonferroni correction (family-wise $\alpha=0.05$; $m=6$, per-test $\alpha=0.0083$) in pairwise comparisons between $F_{6}$ and $f_{1},\cdots,f_{6}$.
IV-C Results on Calibration to the Synthetic Data
Note that although the non-identifiable set of any $F_{i}$ is verified to be smaller than that of any $F_{j}$, $i>j$, it does not deterministically lead to better calibration performance of $F_{i}$ than $F_{j}$. Because $F$ mainly decides the parameter space, while the calibration performance is also jointly influenced by the optimization algorithms. Given the calibration problem is black-box with multiple local optima, it is non-trivial for the optimization algorithms to find the global optimum even if the parameter space is modeled fully identifiable.
The performance between each simulated data and its corresponding synthetic data is measured by the indicators based on both MSE and W, respectively. In Fig. 8, each (blue or yellow) point displays the average performance on these 10 synthetic data. It shows that by using more features in $F$, the calibration performance generally improves as the distance values decrease.
Furthermore, it is interesting to see how the simulation model calibrated by all features ($F_{6}$) compares to the ones calibrated to every single feature $f_{i},i=1,...,6$, instead of in an incremental manner of $F_{i},i=1,...,5$. As listed in Table III, by calibrating with $F_{6}$, the obtained parameter enjoys the best simulation performance, i.e., the smallest averaged distances on 7 out of the 10 instances and the top 3 averaged distances on the other 3 instances. Besides, the Friedman test ($p=6.7\times 10^{-4}$) indicates significant ranking differences among methods. Subsequent pairwise Wilcoxon signed-rank tests, with Bonferroni correction, confirm that $F_{6}$ consistently outperforms $f_{1}$ to $f_{6}$ with statistical significance. This implies that the proposed theory and new objective function are applicable to various features and not restricted by the feature with more information about the social system, e.g., the mid-price.
We further expand the experiments in Fig. 8 to more different values of $\epsilon$. Note that $\epsilon$ determines how the random noise in the financial market data impacts the measure of the discrepancy between two time series. The higher the value of $\epsilon$ is, the more likely two time series are indistinguishable given Definition  1. The results are listed in Table IV. It can be seen that under different $\epsilon$, the calibration performance always decreases linearly with the increasing number of features. This shows that the proposed objective function is robust with respect to different observation noises.
@@ IV-C Results on Calibration to the Synthetic Data
Then, we illustrate the time series of 6 features to show how the simulated data resembles the synthetic data in Fig. 10. Specifically, we calibrate the model by $F_{6}$ and $f_{x}$ on the data generated based on the recommended parameter setting $\omega^{*}=[0.025,100,10,0.001,0.15,0.025]$ [8], where $f_{x}$ indicates any single feature which helps to calibrate the model and generate the best-simulated data. The simulated data calibrated by $F_{6}$ closely resembles the target data on all 6 features. To our best knowledge, this is the first ABM-agnostic method of successfully calibrating the high-frequency market time series at 1 second level and simulating 6 features with high-fidelity. In comparison, the simulated data of $f_{x}$ performs much poorer. This immediately suggests that calibrating more features can lead to low non-identifiability and high simulation fidelity.
TABLE IV: Linear regression coefficient table.*
$\epsilon$ evaluation $D$ slope intercept $R^{2}$
0.016 W -1.05 9.44 0.61
MSE -0.185 4.18 0.59
0.024 W -1.11 9.28 0.70
MSE -0.294 4.66 0.84
0.028 W -0.746 8.64 0.54
MSE -0.154 4.18 0.39
0.030 W -0.570 8.70 0.61
MSE -0.167 4.48 0.24
    Note that the unit of slope and intercept is at the $1E10^{-2}$ level.
IV-D Ablation on Aggregation and Discrepancy
The proof in Appendix A-C shows that maximization uniquely preserves the intersection property of non-identifiable sets. To empirically validate this, Table V presents an ablation study comparing aggregation methods (max, min, mean) and calibration measures (KL divergence (KL), Kolmogorov–Smirnov test (KS), MSE, W) across 10 synthetic datasets. The Friedman test confirms significant performance differences ($p=2.88\times 10^{-3}$). Subsequent Wilcoxon signed-rank tests show that max-W significantly outperforms min-W ($p=4.88\times 10^{-3}$) and mean-W ($p=9.8\times 10^{-4}$), empirically validating the theoretical claim that maximization enhances identifiability. Within the max group, max-W achieves the lowest rank (1.50), significantly outperforming KL ($p=4.88\times 10^{-3}$), KS ($p=6.84\times 10^{-3}$), and MSE ($p=0.0527$), making W the preferable choice for FMS.
TABLE V: Statistical ablation tests of aggregation methods and calibration measures*.
aggregate max min mean
measure KL KS MSE W W W
data 1 3.19 4.47 3.47 3.07 6.63 4.77
data 2 6.03 3.08 3.03 2.90 3.36 3.89
data 3 6.86 4.30 3.19 3.08 4.63 3.47
data 4 3.01 3.10 3.90 2.47 4.32 4.33
data 5 2.55 6.58 3.75 2.84 3.01 3.79
data 6 3.71 6.69 4.76 3.67 3.74 5.73
data 7 3.93 3.10 1.42 3.51 3.23 3.52
data 8 3.82 3.18 4.65 2.26 3.88 2.54
data 9 8.35 5.48 3.85 3.51 5.07 5.40
data 10 3.69 3.63 2.10 2.23 5.32 4.04
#rank $3.90^{+}$ $3.90^{+}$ $2.90^{-}$ 1.50 $4.30^{+}$ $4.50^{+}$
    Performance is at the $1E10^{-2}$ level. $+$ denotes significance under Bonferroni correction (family-wise $\alpha=0.05$; $m=5$, per-test $\alpha=0.01$) in pairwise comparisons between max-W and others, and - indicates no significance.
@@ IV-D Ablation on Aggregation and Discrepancy
IV-E Results on Calibration to the Real Data
At last, it is verified that the proposed method can also calibrate the real market data at a 3-second level. A 1-hour time series of 000001.sz is calibrated using $F_{1}$ and $F_{3}$. The simulated data generated by the corresponding best calibrated models is depicted in Fig. 11, together with the real observed data. Note that most of the FMS simulators (including the adopted PGPS model) model the order volume in a coarse-grained manner, i.e., simply setting all the orders with the identical volume of 100. Hence, only the mid-price time series is shown in a time step of 1 minute. It can be seen that the simulated mid-price of $F_{3}$ is more similar to the real observed data than $F_{1}$ by an improvement of 37.9% in MSE. This confirms that the proposed method indeed can calibrate the real market data with multiple time series, thus suggesting a new effective calibration objective function for FMS.
V Conclusions and Future Work
@@
This paper studies the non-identifiability issue of social simulation. To our best knowledge, this is the first work in which the non-identifiability is formally defined and theoretically alleviated. It is defined as the probability of random sampling in the non-identifiable set of the parameter space. It is rigorously analyzed that this probability can be reduced with more time series features in an exponential manner. Based on the analysis, a new objective function is proposed to effectively aggregate multiple features with a maximization function over existing discrepancy metrics. We theoretically and empirically show that the proposed aggregation is the optimal one. Extensive empirical studies have been conducted on 10 synthetic data and 1 real data of the FMS tasks. It has been successfully verified through simulations that, by using multiple features, not only can the non-identifiable set be minimized, but the simulation fidelity is significantly improved. Finally, we suggest how to select features using theoretical and empirical studies.
Future work can be multi-directional. First, while the study relies on handcrafted LOB features, learning-based methods, such as advanced representation learning techniques [21] and interpretable alpha factor discovery [36], could help extract more meaningful market signals and refine feature selection. Second, developing adaptive recalibration methods that dynamically adjust parameters based on incoming market data could help maintain high-fidelity simulations over longer time horizons and under changing conditions. Third, beyond financial markets, the approach of reducing non-identifiability through multivariate feature calibration could extend to other agent-based simulations that generate multivariate time series data, like weather forecasting and traffic flow modeling.
Appendix A Appendix
A-A Proof of inequality (7c)
Proof 1
Recall that the inequality is
$\prod_{i=1}^{K}\beta_{i}\leq\left(\frac{\sum_{i=1}^{K}\beta_{i}^{2}}{K}\right)^{\frac{K}{2}}.$
$\forall i\in\{1,\cdots,K\},$ we have $\beta_{i}>0$; by definition, $\beta_{i}\leq 1$.
Consider $f(x)=\log_{e}^{x}\triangleq\ln x$, which is concave on $(0,+\infty)$. Suppose $p_{i}\geq 0$ and $\sum_{i=1}^{K}p_{i}=1$, by Jensen’s inequality, we have
$p_{i}\sum_{i=1}^{K}\ln x_{i}\leq\ln\left(\sum_{i=1}^{K}p_{i}x_{i}\right).$ (11)
Replace $p_{i}$ and $x_{i}$ with $\frac{1}{K}$ and $\beta_{i}^{2}$, respectively, then
$\frac{1}{K}\sum_{i=1}^{K}\ln\beta_{i}^{2}\leq\ln\left(\frac{1}{K}\sum_{i=1}^{K}\beta_{i}^{2}\right).$ (12)
Also
$\frac{1}{K}\sum_{i=1}^{K}\ln\beta_{i}^{2}=\sum_{i=1}^{K}\ln\beta_{i}^{\frac{2}{K}}=\ln\left(\ \left(\prod_{i=1}^{K}\beta_{i}\right)^{\frac{2}{K}}\right).$ (13)
By combining (12) with (13), and noticing the fact that $\ln x$ is strictly increasing on $(0,+\infty)$, we get the desired result. $\blacksquare$
A-B Proof of Theorem 1.
Proof 2
To prove the exponential alleviation of the upper bound of non-identifiability under the constraint
$\beta_{K}^{2}<\frac{{\sum}_{i=1}^{K-1}\beta_{i}^{2}}{K-1},$ (14)
we consider two cases with $K-1$ and $K$ features, respectively. The non-identifiability bounds in these cases are given by
$P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K-1}\middle|\hat{\mathbf{X}}_{T}\right)\leq\left(\frac{1}{K-1}\sum_{i=1}^{K-1}\beta_{i}^{2}\right)^{\frac{K-1}{2}}$
and
$P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}\middle|\hat{\mathbf{X}}_{T}\right)\leq\left(\frac{1}{K}\sum_{i=1}^{K}\beta_{i}^{2}\right)^{\frac{K}{2}}.$
We then examine the ratio of these two bounds:
$\frac{\left(\frac{1}{K}\sum_{i=1}^{K}\beta_{i}^{2}\right)^{\frac{K}{2}}}{\left(\frac{1}{K-1}\sum_{i=1}^{K-1}\beta_{i}^{2}\right)^{\frac{K-1}{2}}}=$ (15)
$\left(\frac{K-1}{K}\cdot\left(1+\frac{\beta_{K}^{2}}{\sum_{i=1}^{K-1}\beta_{i}^{2}}\right)\right)^{\frac{K-1}{2}}\cdot\left(\frac{\sum_{i=1}^{K}\beta_{i}^{2}}{K}\right)^{\frac{1}{2}}.$
Given that $\beta_{i}\in(0,1]$ for $i=1,\cdots,K-1$, we have
$0<\left(\frac{\sum_{i=1}^{K}\beta_{i}^{2}}{K}\right)^{\frac{1}{2}}\leq 1$ (16)
Furthermore, from (14), it follows that
$0<\frac{K-1}{K}\cdot\frac{\beta_{K}^{2}}{\sum_{i=1}^{K}\beta_{i}^{2}}<\frac{1}{K}$ (17)
and therefore
$10<\frac{K-1}{K}\cdot\left(1+\frac{\beta_{K}^{2}}{\sum_{i=1}^{K-1}\beta_{i}^{2}}\right)<$ 1 . (18)
Define
$\Delta_{K}=1-\frac{K-1}{K}\cdot\left(1+\frac{\beta_{K}^{2}}{\sum_{i=1}^{K-1}\beta_{i}^{2}}\right),$ (19)
then $10<\Delta_{K}<$ 1. Substituting into the $RHS$ of (15) yields
[…]
Let $\lambda_{K}\triangleq-\frac{1}{2}\ln(1-\Delta_{K})>0$. Then $(1-\Delta_{K})^{\frac{K}{2}}=e^{-\lambda_{K}K}$. Since $\inf_{K}\Delta_{K}>0$, it follows that $\inf_{K}\lambda_{K}>0$. Define the uniform constant $\lambda=\inf_{K}\lambda_{K}>0$, so that $\lambda_{K}\geq\lambda$ for all $K$, and hence $e^{-\lambda_{K}K}\leq e^{-\lambda K}$. Multiplying both sides of (20) by the LHS denominator in (15) (which is less than $1$), and applying the above bound, we obtain
$\left(\frac{1}{K}\sum_{i=1}^{K}\beta_{i}^{2}\right)^{\frac{K}{2}}<\left(1-\Delta_{K}\right)^{\frac{K}{2}}\leq e^{-\lambda\cdot K}.$ (21)
$\blacksquare$
A-C Proof of Uniqueness of Max Aggregation
Proof 3
We consider a common domain $\Omega^{D}$ for all $\Omega^{D,k}$, as originally stipulated in the definition 3.
Define $F_{\max}\{D_{k}\}=\max_{1\leq k\leq K}D_{k}$, where
$D_{k}(\omega)\triangleq D(\hat{\mathbf{X}}_{T}^{k},M(\omega))$ (22)
for clarity of notation. We first establish the utility of $F$, followed by its uniqueness.
If $\omega\in\mathcal{S}^{F_{\max},\epsilon}$, then:
$\max_{k}D_{k}(\omega)\leq\epsilon\implies D_{k}(\omega)\leq\epsilon,\ \forall k\implies$ (23)
$\omega\in\bigcap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}\implies\mathcal{S}^{F_{\max}}\subseteq\bigcap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}$
If $\omega\in\bigcap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}$, then:
$D_{k}(\omega)\leq\epsilon,\ \forall k\implies\max_{k}D_{k}(\omega)\leq\epsilon\implies$ (24)
$\omega\in\mathcal{S}^{F_{\max},\epsilon}\implies\bigcap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}\subseteq\mathcal{S}^{F_{\max},\epsilon}$
Combine the results, we have
$\mathcal{S}^{F_{\max},\epsilon}=\bigcap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon},\quad\forall\epsilon>0,\epsilon\in\mathbb{R}$ (25)
which completes the proof of the utility of $F$. Next, we establish the uniqueness of $F$ via contradiction. Suppose these exists $F^{\prime}\neq F_{\max}$ that satisfies (25). Consider two cases:
$\exists\{D_{k}^{*}\}$ s.t. $F^{\prime}(\{D_{k}^{*}\})<\max_{k}D_{k}^{*}$. Take $\epsilon_{0}=F^{\prime}(\{D_{k}^{*}\})$:
$\max_{k}D_{k}^{*}>\epsilon_{0}\ \text{but}\ \omega\in\mathcal{S}^{F^{\prime},\epsilon_{0}}$ (26)
$\omega\in\mathcal{S}^{F^{\prime},\epsilon_{0}}\implies\omega\in\bigcap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon_{0}}\implies D_{k}^{*}\leq\epsilon_{0},\ \forall k.$
Contradicts $\max_{k}D_{k}^{*}>\epsilon_{0}$.
$\exists\{D_{k}^{*}\}$ s.t. $F^{\prime}(\{D_{k}^{*}\})>\max_{k}D_{k}^{*}$. Take $\epsilon_{0}=\max_{k}D_{k}^{*}$:
$\omega\in\bigcap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon_{0}}\implies D_{k}^{*}\leq\epsilon_{0},\ \forall k\implies F^{\prime}(\{D_{k}^{*}\})\leq\epsilon_{0}.$ (27)
Contradicts $F^{\prime}(\{D_{k}^{*}\})>\epsilon_{0}$, thus completing the proof and establishing that maximization is the unique aggregation method that preserves the intersection of non-identifiable sets.
Moreover, we analyze two common alternative aggregation methods. The minimum aggregation function is defined as $F_{\min}\{D_{k}\}=\min_{1\leq k\leq K}D_{k}$, This leads to
$\mathcal{S}^{F_{\min},\epsilon}=\bigcup_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}.$ (28)
The weighted sum aggregation function is defined as $F_{sum}\{D_{k}\}=\sum_{k=1}^{K}\eta_{k}D_{k}$, where $\eta_{k}>0$ and $\sum_{k=1}^{K}\eta_{k}=1$. This results in
$\mathcal{S}^{F_{sum},\epsilon}\subseteq\bigcap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon/\eta_{k}}.$ (29)
$\blacksquare$
A-D Proof of Selecting Linearly Independent Features
Proof 4
We demonstrate that if two features are correlated by an affine transformation, there exists a distance metric for which their non-identifiable sets are identical. Therefore, features should be chosen to be linearly independent.
Let $\hat{\mathbf{X}}_{T}^{k}$ and $\mathbf{X}_{T}^{k}$ (for $k=1,2$) represent target and generated time series features related by:
$\begin{cases}\hat{\mathbf{X}}_{T}^{1}=a\hat{\mathbf{X}}_{T}^{2}+b\\ \mathbf{X}_{T}^{1}=a\mathbf{X}_{T}^{2}+b\end{cases},\quad a\neq 0,\ b\in\mathbb{R}.$ (30)
The non-identifiable sets are:
$S^{D,k,\epsilon}=\left\{\omega\middle|D\left(\hat{\mathbf{X}}_{T}^{k}(\omega),\mathbf{X}_{T}^{k}\right)\leq\epsilon\right\}.$ (31)
Choose $D$ as the standardized MSE distance, denoted as $D_{\text{std}}(X,Y)$,
$D_{\text{std}}(\mathbf{X}_{T}^{k},\mathbf{Y}_{T}^{k})\triangleq\frac{1}{T}\sum_{t=t_{s}}^{t_{e}}\left(\frac{X_{t}-\mu_{\mathbf{X}_{T}^{k}}}{\sigma_{\mathbf{X}_{T}^{k}}}-\frac{Y_{t}-\mu_{\mathbf{Y}_{T}^{k}}}{\sigma_{\mathbf{Y}_{T}^{k}}}\right)^{2}$ (32)
and the mean and standard deviation is defined as
$\mu_{\mathbf{X}_{T}^{k}}\triangleq\frac{1}{T}\sum_{t=t_{s}}^{t_{e}}X_{t}$ (33)
$\sigma_{\mathbf{X}_{T}^{k}}\triangleq\sqrt{\frac{1}{T}\sum_{t=t_{s}}^{t_{e}}(X_{t}-\mu_{\mathbf{X}_{T}^{k}})^{2}}.$
For affine-related features, we have
$\frac{\mathbf{X}_{T}^{1}-\mu_{\mathbf{X}_{T}^{1}}}{\sigma_{\mathbf{X}_{T}^{1}}}=\text{sign}(a)\frac{\mathbf{X}_{T}^{2}-\mu_{\mathbf{X}_{T}^{2}}}{\sigma_{\mathbf{X}_{T}^{2}}}.$ (34)
the relationship for $\hat{\mathbf{X}}_{T}^{2}$ and $\mathbf{X}_{T}^{2}$ is analogous, which yields
$D_{\text{std}}\left(\hat{\mathbf{X}}_{T}^{1},\mathbf{X}_{T}^{1}\right)=D_{\text{std}}\left(\hat{\mathbf{X}}_{T}^{2},\mathbf{X}_{T}^{2}\right).$ (35)
Consequently, the non-identifiable sets are identical, i.e.,
$S^{D_{\text{std}},1,\epsilon}=S^{D_{\text{std}},2,\epsilon}.$ (36)
Under the max-aggregation $F$, we obtain
$S^{F,\epsilon}=S^{D_{\text{std}},1,\epsilon}\cap S^{D_{\text{std}},2,\epsilon}=S^{D_{\text{std}},k,\epsilon},\quad\forall k.$ (37)
Thus, the measure remains unchanged:
$\mu(S^{F,\epsilon})=\mu(S^{D_{\text{std}},k,\epsilon})\quad\forall k.$ (38)
$\blacksquare$